OpenStudy (anonymous):

-1=5+x/6 I'm trying to help my son, but I can't figure this out. Help!

OpenStudy (anonymous):

-1-5 = x/6 -6 = x/6 -36= x

OpenStudy (anonymous):

According to his teacher, she have the answer as x= -11, he has to figure out how. I got the above answer as well

OpenStudy (radar):

-6=5+x x=-11